OpenStudy (anonymous):

how do i do this equation? square root of 20x^3

OpenStudy (anonymous):

you mean "write in simplest radical form"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{20x^3}=\sqrt{4\times 5\times x^2\times x}=\sqrt{4}\sqrt{x^2}\sqrt{5x}=2x\sqrt{5x}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

mind giving me an idea of this one? its a little different. 3(in the index) square root of 81x^2